OpenStudy (anonymous):

Find the exact value of the trigonometric function at the given real number? - Sin (21pi/2) - Cos (21pi/2) - Cot (21pi/2)

OpenStudy (anonymous):

Sin(21pi/2)=1 Cos(21pi/2)=0 Cot(21pi/2)=cos(21pi/2)/sin(21pi/2)=0/1=0